# Stochastic Gradient Langevin Unlearning

### v1

@@ Abstract
“The right to be forgotten” ensured by laws for user data privacy becomes increasingly important. Machine unlearning aims to efficiently remove the effect of certain data points on the trained model parameters so that it can be approximately the same as if one retrains the model from scratch. This work proposes stochastic gradient Langevin unlearning, the first unlearning framework based on noisy stochastic gradient descent (SGD) with privacy guarantees for approximate unlearning problems under convexity assumption. Our results show that mini-batch gradient updates provide a superior privacy-complexity trade-off compared to the full-batch counterpart. There are numerous algorithmic benefits of our unlearning approach, including complexity saving compared to retraining, and supporting sequential and batch unlearning. To examine the privacy-utility-complexity trade-off of our method, we conduct experiments on benchmark datasets compared against prior works. Our approach achieves a similar utility under the same privacy constraint while using $2\%$ and $10\%$ of the gradient computations compared with the state-of-the-art gradient-based approximate unlearning methods for mini-batch and full-batch settings, respectively.
Machine unlearning, privacy, stochastic gradient descent, stochastic gradient Langevin dynamic
@@ 1 Introduction
Machine learning models are now essential tools used in industry with applications ranging from biology, computer vision, to natural language processing. Data that these models learn from are often collected from users by corporations for which user data privacy has to be respected. Certain laws, including the European Union’s General Data Protection Regulation (GDPR), are in place to ensure “the right to be forgotten”, which requires corporations to erase all information pertaining to a user if they request to remove their data. It is insufficient to comply with such privacy regulation by only removing user data from the dataset, as machine learning models are able to memorize training data information and risk information leakage (Carlini et al., 2019; Guo et al., 2023). Therefore, it is also critical to erase such information from the machine learning model. This problem is now an important research direction known as machine unlearning (Cao & Yang, 2015).
A naive approach to achieving machine unlearning is to retrain the model from scratch after every data removal request, which leads to a “perfect” privacy guarantee. Apparently, this approach is prohibitively expensive in practice for frequent data removal requests. Various machine unlearning strategies have been proposed to avoid retraining from scratch with a better complexity. They can be roughly categorized into exact (Bourtoule et al., 2021; Ullah et al., 2021; Ullah & Arora, 2023) as well as approximate approaches (Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Chien et al., 2022, 2024). Exact approaches ensure that the unlearned model would be identical to the retraining one in distribution. In contrast, a slight misalignment between the unlearned model and the retraining one in distribution is allowed for approximate approaches with a notion similar to Differential Privacy (DP) (Dwork et al., 2006). Interestingly, this probabilistic notion of approximate unlearning naturally leads to a privacy-utility trade-off for machine unlearning.
Figure 1: The overview of our stochastic gradient Langevin unlearning framework. (Left) Given a mini-batch sequence $\mathcal{B}$, the learning process $\mathcal{M}$ induces a regular polyhedron where each vertex corresponds to a stationary distribution $\nu_{\mathcal{D}|\mathcal{B}}$ for each dataset $\mathcal{D}$. $\nu_{\mathcal{D}|\mathcal{B}}$ and $\nu_{\mathcal{D|\mathcal{B}}^{\prime}}$ are adjacent if $\mathcal{D},\mathcal{D}^{\prime}$ differ in one data point. There is an uniform upper bound $Z_{\mathcal{B}}$ for the infinite Wasserstein distance $W_{\infty}$ between any pair of adjacent distributions $\nu_{\mathcal{D}|\mathcal{B}}$ and $\nu_{\mathcal{D|\mathcal{B}}^{\prime}}$. Intuitively, smaller $Z_{\mathcal{B}}$ implies an “simpler” unlearning problem, since transporting from $\nu_{\mathcal{D}|\mathcal{B}}$ to $\nu_{\mathcal{D|\mathcal{B}}^{\prime}}$ is easier. (Right) Learning and unlearning processes on adjacent datasets. We apply PABI analysis to convert the initial $W_{\infty}$ bound to Rényi difference bound $d_{\alpha}^{\mathcal{B}}$, and apply joint convexity of KL divergence to obtain the final privacy loss $\varepsilon$, which also take the randomness of $\mathcal{B}$ into account.
1.1 Our Contributions
We propose stochastic gradient Langevin unlearning, an approximate unlearning framework based on projected noisy stochastic gradient descent (PNSGD). To the best of our knowledge, we are the first to provide a thorough analysis of mini-batch gradient methods for approximate unlearning problems. Mini-batch private learning algorithms, such as DP-SGD (Abadi et al., 2016), have achieved great success due to their scalability compared to their full-batch counterparts. Prior works (Neel et al., 2021; Chien et al., 2024) focus on full-batch gradient algorithms while the benefit of the mini-batch gradient is still unclear for approximate unlearning problems. Our results show that PNSGD can be well applied to approximate unlearning problems with a substantial improvement in privacy-utility-complexity trade-offs for practical usage. Moreover, our stochastic gradient Langevin unlearning framework brings various benefits, including (1) provable computational benefit compared to retraining from scratch and (2) naturally supporting sequential and batch unlearning for multiple data point removal.
@@ 1.1 Our Contributions
The core idea of our stochastic gradient Langevin unlearning is illustrated in Figure 1. Given a training dataset $\mathcal{D}$ and a fixed mini-batch sequence $\mathcal{B}$, we learn the model parameter with PNSGD. For sufficient learning epochs, we prove that the law of the PNSGD learning process converges to an unique stationary distribution $\nu_{\mathcal{D}|\mathcal{B}}$ (Theorem 3.1). When an unlearning request arrives, we update $\mathcal{D}$ to an adjacent dataset $\mathcal{D}^{\prime}$ so that the data point subject to such request is removed. The approximate unlearning problem can then be viewed as moving from the current distribution $\nu_{\mathcal{D}|\mathcal{B}}$ to the target distribution $\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$ until $\varepsilon$-close in Rényi divergence for the desired privacy loss $\varepsilon$ [^1] We refer privacy loss as two-sided Rényi divergence of two distributions, which we defined as Rényi difference in Definition 2.1.. Intuitively, if the starting distribution $\nu_{\mathcal{D}|\mathcal{B}}$ and the target distribution $\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$ are “close”, we require less unlearning epochs to achieve a desired $\varepsilon$.
We prove this intuition formally in Theorem 3.2. We leverage the Privacy Amplification By Iteration (PABI) analysis (Feldman et al., 2018; Altschuler & Talwar, 2022a, b), which allows us to convert an initial infinite Wasserstein ($W_{\infty}$) distance to a Rényi divergence bound. We show that $K$ unlearning epochs lead to an exponentially fast decay in privacy loss from the initial $W_{\infty}$ distance bound $Z_{\mathcal{B}}$ of any pair of adjacent stationary distributions $\nu_{\mathcal{D}|\mathcal{B}},\nu_{\mathcal{D|\mathcal{B}}^{\prime}}$ under the strong convexity assumption. We later generalize it to the convex-only case (Corollary 3.10). More specifically, we show that $K$ unlearning epochs lead to privacy loss $\varepsilon=O(Z_{\mathcal{B}}^{2}c^{2Kn/b})$ for mini-batch size $b$, where we also characterize $Z_{\mathcal{B}}$ by tracking the $W_{\infty}$ distance of any two learning processes on adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. To the best of our knowledge, we are the first to introduce PABI analysis for unlearning problems and achieve approximate unlearning guarantees for PNSGD. To prevent the worst-case bound induced by the fixed mini-batch sequence $\mathcal{B}$, we further utilize the randomness in the selection of $\mathcal{B}$ for an improved privacy bound (Corollary 3.9). Our approach naturally extends to multiple unlearning requests, including sequential and batch unlearning settings, which can be done by extending our analysis of tracking the $W_{\infty}$ distance bound $Z_{\mathcal{B}}$ for these cases (Theorem 3.11 and Corollary 3.12).
Our results demonstrate that smaller batch size $b$ leads to a better privacy loss decaying rate due to the more updates per epoch. Intuitively, more unlearning updates imply closer to the “retraining distribution” $\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$ and thus better privacy. It highlights the benefit of the mini-batch setting for unlearning compared to prior works that only focus on full-batch gradient methods. Nevertheless, an extremely small $b$ can degrade the model utility and incur instability in practice, as the stationary distribution $\nu_{\mathcal{D}|\mathcal{B}}$ also depends on the design of mini-batches $\mathcal{B}$. Our results highlight a unique privacy-utility-complexity trade-off with respect to the mini-batch size $b$ for approximate unlearning.
Regarding the computational benefit compared to retraining from scratch, we may show epoch saving by comparing two $W_{\infty}$ distances, the one between initialization $\nu_{0}$ and $\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$, which is $\Omega(1)$ in the worst case versus the other one between the adjacent distributions $\nu_{\mathcal{D}|\mathcal{B}}$ and $\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$, which is upper bounded by $Z_{\mathcal{B}}\approx O(\eta M/b)$ for bounded gradient norm $M$ and step size $\eta$. For moderate mini-batch size $b$, our unlearning approach is more efficient than retraining from scratch. Even if one chooses $b=1$ (which leads to worse utility as mentioned earlier), our approach can still be more efficient as long as $\eta M$ is properly controlled. Our experiment shows that commonly used mini-batch size ($b\geq 32$) already leads to significant computational savings compared to retraining.
@@
To examine the trade-off between privacy-utility-complexity of stochastic gradient Langevin unlearning, we conduct experiments through logistic regression tasks on benchmark datasets following a similar setting of Guo et al. (2020). We compared with the state-of-the-art gradient-based approximate unlearning solutions (Neel et al., 2021; Chien et al., 2024), albeit they only focus on full-batch setting $b=n$. Our analysis provides a significantly better privacy-utility-complexity trade-off even when we restrict ourselves to $b=n$, and further improves the results by adopting mini-batches. Under the same privacy constraint, stochastic gradient Langevin unlearning achieves similar utility while merely requiring $10\%,2\%$ of gradient computations compared to baselines for full and mini-batch settings respectively. It demonstrates the benefit of using mini-batches for approximate unlearning.
@@ 1.2 Related Works
Machine unlearning with privacy guarantees. The concept of approximate unlearning uses a probabilistic definition of $(\epsilon,\delta)$-unlearning motivated by differential privacy (Dwork et al., 2006), which is studied by Guo et al. (2020); Sekhari et al. (2021); Chien et al. (2022). Notably, the unlearning approach of these works involved Hessian inverse computation which can be computationally prohibitive in practice for high dimensional problems. Ullah et al. (2021) focus on exact unlearning via a sophisticated version of noisy SGD. Their analysis is based on total variation stability which is not directly applicable to approximate unlearning settings and different from our analysis focusing on Rényi divergence.  Neel et al. (2021) leverage full-batch PGD for (un)learning and achieve approximate unlearning by publishing the final parameters with additive Gaussian noise.  Gupta et al. (2021); Ullah & Arora (2023); Chourasia & Shah (2023) study the setting of adaptive unlearning requests, where the unlearning request may depend on the previous (un)learning results. It is possible to show that our framework is also capable of this adaptive setting since we do not keep any non-private internal states, albeit we only focus on non-adaptive settings in this work. We left a rigorous discussion for this as future work.
Comparison to Chien et al. (2024). Very recently, Chien et al. (2024) propose to leverage PNGD for learning and unlearning, where they establish approximate unlearning guarantees based on Langevin dynamic analaysis (Vempala & Wibisono, 2019) and track the Log-Sobolev conditions along the (un)learning processes (Gross, 1975; Chewi, Sinho, 2023). In contrast, we focus on the mini-batch setting (PNSGD), where our analysis is based on PABI and we track the $W_{\infty}$ distance along the (un)learning processes. There are pros and cons to these two types of analyses. On one hand, our results based on PABI analysis provide a better privacy-utility-complexity trade-off for sequential unlearning under a strongly convex setting. More specifically, we leverage standard triangle inequality for analyzing $W_{\infty}$ distance while Chien et al. (2024) require to adopt weak triangle inequality of Rényi divergence for each sequential unlearning request. This results in a worse estimation of sequential unlearning. On the other hand, Langevin dynamic analysis gives a unified analysis for non-convex settings where the convexity assumption of PABI analysis cannot be dropped (Altschuler & Talwar, 2022a, b). It is interesting to further explore the limit of these two analyses for unlearning problems, which we left as future work.
@@ 2 Preliminaries
Definition 2.2 (Rényi Differential Privacy (RDP) (Mironov, 2017)).
@@
It is known to the literature that an $(\alpha,\varepsilon)$-RDP guarantee can be converted to the popular $(\epsilon,\delta)$-DP guarantee (Dwork et al., 2006) relatively tight (Mironov, 2017). As a result, we will focus on establishing results with respect to $\alpha$ Rényi difference (and equivalently $\alpha$ Rényi difference). Next, we introduce our formal definition of unlearning based on $\alpha$ Rényi difference as well.
@@
Consider a randomized learning algorithm $\mathcal{M}:\mathcal{X}^{n}\mapsto\mathbb{R}^{d}$ and a randomized unlearning algorithm $\mathcal{U}:\mathcal{R}^{d}\times\mathcal{X}^{n}\times\mathcal{X}^{n}\mapsto\mathcal{R}^{d}$. We say $(\mathcal{M},\mathcal{U})$ achieves $(\alpha,\varepsilon)$-RU if for any $\alpha>1$ and any adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$, the $\alpha$ Rényi difference $d_{\alpha}(\rho,\nu^{\prime})\leq\varepsilon,$ where $\mathcal{U}(\mathcal{M}(\mathcal{D}),\mathcal{D},\mathcal{D}^{\prime})\sim\rho$ and $\mathcal{M}(\mathcal{D}^{\prime})\sim\nu^{\prime}$.
Our Definition 2.3 can be converted to the standard $(\epsilon,\delta)$-unlearning definition defined in Guo et al. (2020); Sekhari et al. (2021); Neel et al. (2021), similar to RDP to DP conversion (see Appendix K). Since we work with the replacement definition of dataset adjacency, to unlearn a data point $\mathbf{d}_{i}$ we can simply replace it with any data point $\mathbf{d}_{i}^{\prime}\in\mathcal{X}$ for the updated dataset $\mathcal{D}^{\prime}$ in practice. One may also repeat the entire analysis with the objective function being the summation of individual loss for the standard add/remove notion of dataset adjacency. Finally, we will also leverage the infinite Wasserstein distance in our analysis.
@@
Algorithm 1 Stochastic gradient Langevin unlearning
@@
1:  Minibatch Generation: randomly partition indices $[n]$ into $n/b$ mini-batches of size $b$: $\mathcal{B}^{0},\ldots,\mathcal{B}^{n/b-1}$.
@@
2:  Learning process $\mathcal{M}(\mathcal{D})$: given a dataset $\mathcal{D}=\{\mathbf{d}_{i}\}_{i=1}^{n}\in\mathcal{X}^{n}$, sample initial parameter $x_{0}^{0}$ from a given initialization distribution $\nu_{0}$ supported on $\mathcal{C}_{R}$.
3:  for  epoch $t=0,\ldots,T-1$ do
4:     for  iteration $j=0,\ldots,n/b-1$ do
@@
5:        $x^{j+1}_{t}=\Pi_{\mathcal{C}_{R}}\left(x^{j}_{t}-\eta g(x^{j}_{t},\mathcal{B}^{j})+\sqrt{2\eta\sigma^{2}}W^{j}_{t}\right),$ where $g(x^{j}_{t},\mathcal{B}^{j})=\frac{1}{b}\sum_{i\in\mathcal{B}^{j}}\nabla f(x^{j}_{t};\mathbf{d}_{i})$ and $W^{j}_{t}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(0,I_{d})$.
6:     end for
7:     $x^{0}_{t+1}=x^{n/b}_{t}$
8:  end for
@@
9:  Unlearning process $\mathcal{U}(\mathcal{M}(\mathcal{D}),\mathcal{D}^{\prime})$: given an updated dataset $\mathcal{D}^{\prime}=\{\mathbf{d}_{i}^{\prime}\}_{i=1}^{n}\in\mathcal{X}^{n}$ and current parameter $y_{0}^{0}$,
10:  for  epoch $k=0,\ldots,K-1$ do
11:     for  iteration $j=0,\ldots,n/b-1$ do
@@
12:        $y^{j+1}_{k}=\Pi_{\mathcal{C}_{R}}\left(y^{j}_{k}-\eta g(y^{j}_{k},\mathcal{B}^{j})+\sqrt{2\eta\sigma^{2}}W^{j}_{k}\right),$ where $g(y^{j}_{k},\mathcal{B}^{j})=\frac{1}{b}\sum_{i\in\mathcal{B}^{j}}\nabla f(y^{j}_{k};\mathbf{d}_{i}^{\prime})$ and $W^{j}_{k}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(0,I_{d})$.
13:     end for
14:     $y^{0}_{k+1}=y^{n/b}_{k}$
15:  end for
3 Stochactic Gradient Langevin Unlearning
We now introduce our stochastic gradient Langevin unlearning strategy. We propose to leverage the projected noisy SGD (PNSGD) with a cyclic mini-batch strategy. Note that this mini-batch strategy is not only commonly used for practical DP-SGD implementations in privacy libraries (Yousefpour et al., 2021), but also in theoretical analysis for DP guarantees (Ye & Shokri, 2022). For the learning algorithm $\mathcal{M}$, we propose to optimize the objective function with projected noisy SGD on dataset $\mathcal{D}$ (line 3-8 in Algorithm 1). $\eta,\sigma^{2}>0$ are hyperparameters of step size and noise variance respectively. The initialization $\nu_{0}$ is an arbitrary distribution supported on $\mathcal{C}_{R}$ if not specified. For the unlearning algorithm $\mathcal{U}$, we fine-tune the current parameter on the updated dataset $\mathcal{D}^{\prime}$ subject to the unlearning request (line 10-15 in Algorithm 1) with $y_{0}^{0}=x_{T}^{0}=\mathcal{M}(\mathcal{D})$. For the rest of the paper, we denote $\nu_{t}^{j},\rho_{k}^{j}$ as the probability density of $x_{t}^{j},y_{k}^{j}$ respectively. Furthermore, we denote =$\mathcal{B}=\{\mathcal{B}^{j}\}_{j=0}^{n/b-1}$ the minibatch sequence described in Algorithm 1, where $b$ is the step size and we assume $n$ is divided by $b$ throughout the paper for simplicity [^2] When $n$ is not divided by $b$, we can simply drop the last $n-\lfloor n/b\rfloor b$ points.. We use $\nu_{\cdot|\mathcal{B}}$ to denote the conditional distribution of $\nu_{\cdot}$ given $\mathcal{B}$.
3.1 Limiting distribution and general idea
Now we introduce the general idea of our stochastic gradient Langevin unlearning, which is illustrated in Figure 1. We first prove that for any fixed mini-batch sequence $\mathcal{B}$, the limiting distribution $\nu_{\mathcal{D}|\mathcal{B}}$ of the learning process exists, is unique, and stationery.
Theorem 3.1.
Suppose that the closed convex set $\mathcal{C}_{R}\subset\mathbb{R}^{d}$ is bounded with $\mathcal{C}_{R}$ having a positive Lebesgue measure and that $\nabla f(\cdot;\mathbf{d}_{i}):\mathcal{C}_{R}\to\mathbb{R}^{d}$ is continuous for all $i\in[n]$. The Markov chain $\{x_{t}:=x_{t}^{0}\}$ in Algorithm 1 for any fixed mini-batch sequence $\mathcal{B}$ admits a unique invariant probability measure $\nu_{\mathcal{D}|\mathcal{B}}$ on the Borel $\sigma$-algebra of $\mathcal{C}_{R}$. Furthermore, for any $x\in\mathcal{C}_{R}$, the distribution of $x_{t}$ conditioned on $x_{0}=x$ converges weakly to $\nu_{\mathcal{D}|\mathcal{B}}$ as $t\to\infty$.
The proof of Theorem 3.1 is deferred to Appendix C and is based on applying the results in Meyn & Tweedie (2012) to establish the ergodicity of the learning process $x_{t}^{0}$.
3.2 Unlearning guarantees with PABI analysis
We prove the following unlearning guarantee for stochastic gradient Langevin unlearning described in Algorithm 1 for any fixed mini-batch sequence $\mathcal{B}$. We will later provide a slightly better bound by considering the randomness of $\mathcal{B}$.
Theorem 3.2 (RU guarantee of PNSGD unlearning, fixed $\mathcal{B}$).
Our proof is based on the Privacy Amplification By Iteration (PABI) (Feldman et al., 2018), which has also been leveraged for providing the DP guarantee and mixing time of PNSGD by Altschuler & Talwar (2022a, b). To understand the main idea of the proof, we first introduce the notion of contractive noisy iteration.

### v2

# Certified Machine Unlearning via Noisy Stochastic Gradient Descent

@@ Abstract
“The right to be forgotten” ensured by laws for user data privacy becomes increasingly important. Machine unlearning aims to efficiently remove the effect of certain data points on the trained model parameters so that it can be approximately the same as if one retrains the model from scratch. We propose to leverage projected noisy stochastic gradient descent for unlearning and establish its first approximate unlearning guarantee under the convexity assumption. Our approach exhibits several benefits, including provable complexity saving compared to retraining, and supporting sequential and batch unlearning. Both of these benefits are closely related to our new results on the infinite Wasserstein distance tracking of the adjacent (un)learning processes. Extensive experiments show that our approach achieves a similar utility under the same privacy constraint while using $2\%$ and $10\%$ of the gradient computations compared with the state-of-the-art gradient-based approximate unlearning methods for mini-batch and full-batch settings, respectively.
@@ 1 Introduction
Machine learning models usually learn from user data where data privacy has to be respected. Certain laws, such as European Union’s General Data Protection Regulation (GDPR), are in place to ensure “the right to be forgotten”, which requires corporations to erase all information pertaining to a user if they request to remove their data. It is insufficient to comply with such privacy regulation by only removing user data from the dataset, as machine learning models can memorize training data information and risk information leakage [1, 2]. A naive approach to adhere to this privacy regulation is to retrain the model from scratch after every data removal request. Apparently, this approach is prohibitively expensive in practice for frequent data removal requests and the goal of machine unlearning is to perform efficient model updates so that the resulting model is (approximately) the same as retraining statistically. Various machine unlearning strategies have been proposed, including exact [3, 4, 5, 6] and approximate approaches [7, 8, 9, 10, 11]. The later approaches allow a slight misalignment between the unlearned model and the retraining one in distribution under a notion similar to Differential Privacy (DP) [12].
The most popular approach for privatizing machine learning models with DP guarantee is arguably noisy stochastic gradient methods including the celebrated DP-SGD [13]. Mini-batch training is one of its critical components, which not only benefits privacy through the effect of privacy amplification by subsampling  [14] but also provides improved convergence of the underlying optimization process. Several recent works [9, 11] based on full-batch (noisy) gradient methods may achieve certified approximate unlearning. Unfortunately, their analysis is restricted to the full-batch setting and it is non-trivial to extend these works to the mini-batch setting with tight approximate unlearning guarantees. The main challenge is to incorporate the randomness in the mini-batch sampling into the sensitivity-based analysis [9] or the Langevin-dynamics-based [11] analysis.
@@
We aim to study mini-batch noisy gradient methods for certified approximate unlearning. The high-level idea of our unlearning framework is illustrated in Figure 1. Given a training dataset $\mathcal{D}$ and a fixed mini-batch sequence $\mathcal{B}$, the model first learns and then unlearns given unlearning requests, both via the projected noisy stochastic gradient descent (PNSGD). For sufficient learning epochs, we prove that the law of the PNSGD learning process converges to a unique stationary distribution $\nu_{\mathcal{D}|\mathcal{B}}$ (Theorem 3.1). When an unlearning request arrives, we update $\mathcal{D}$ to an adjacent dataset $\mathcal{D}^{\prime}$ so that the data point subject to such request is removed. The approximate unlearning problem can then be viewed as moving from the current distribution $\nu_{\mathcal{D}|\mathcal{B}}$ to the target distribution $\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$ until $\varepsilon$-close in Rényi divergence for the desired privacy loss $\varepsilon$ [^1] We refer privacy loss as two-sided Rényi divergence of two distributions, which we defined as Rényi difference in Definition 2.1..
Our key observation is that the results of Altschuler and Talwar [15, 16], which study the convergence of PNSGD under the (strong) convexity assumption, can be leveraged after we formulate the approximate unlearning as above. They show that the Rényi divergence of two PNSGD processes with the same dataset but different initial distributions decays at a geometric rate, starting from the infinite Wasserstein distance $(W_{\infty})$ of initial distributions (Figure 1, step 3). As a result, if the initial $W_{\infty}$ distance can be properly characterized, we achieve the corresponding approximate unlearning guarantee by further taking the randomness of the mini-batches $\mathcal{B}$ into account (Figure 1, step 4). Therefore, the key step to establish the PNSGD-based unlearning guarantee is to characterize the initial $W_{\infty}$ distance of the unlearning process tightly.
The projection set diameter $2R$ for this $W_{\infty}$ distance adopted in [15] for DP analysis, unfortunately, leads to a vacuous unlearning guarantee, which cannot illustrate the computational advantage over the retraining from scratch. To alleviate this issue, we perform a careful $W_{\infty}$ distance tracking analysis along the adjacent PNSGD learning processes (Lemma 3.3), which leads to a much better bound $W_{\infty}(\nu_{\mathcal{D}|\mathcal{B}},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})\leq Z_{\mathcal{B}}\approx O(\eta M/b)$ (Figure 1, step 2) for bounded gradient norm $M$, mini-batch size $b$ and step size $\eta$. This ultimately leads to our unlearning guarantee $\varepsilon=O(Z_{\mathcal{B}}^{2}c^{2Kn/b})$ for $K$ unlearning epochs and some rate $c<1$. The computational benefit compared to the retraining from scratch naturally emerges by comparing two $W_{\infty}$ distances, $O(R)$ for the retraining from scratch and $O(\eta M/b)$ for our unlearning framework.
Our approach also naturally extends to multiple unlearning requests, including sequential and batch unlearning settings (Theorem 3.11 and Corollary J.1), by extending the $W_{\infty}$-tracking analysis (Lemma 3.4). Here, we may use the triangle inequality of $W_{\infty}$ distance, which leads to a tigher privacy loss bound (growing linearly to the number of unlearning requests) than that in the Langevin-dynamics-based analysis [11] via weak triangle inequality of Rényi divergence (growing exponentially).
@@
Our results highlight the insights into privacy-utility-complexity trade-off regarding the mini-batch size $b$ for approximate unlearning. A smaller batch size $b$ leads to a better privacy loss decaying rate $O(c^{2Kn/b})$. However, an extremely small $b$ may degrade the model utility and incur instability. It also leads to a worse bound $Z_{\mathcal{B}}\approx O(\eta M/b)$, which degrades the computational benefits compared to retraining. We demonstrate such trade-off of our PNSGD unlearning results via experiments against the state-of-the-art full-batch $(b=n)$ gradient-based approximate unlearning solutions [9, 11]. Our analysis provides a significantly better privacy-utility-complexity trade-off even when we restrict ourselves to $b=n$, and further improves the results by adopting mini-batches. Under the same privacy constraint, our approach achieves similar utility while merely requiring $10\%,2\%$ of gradient computations compared to baselines for full and mini-batch settings respectively.
Figure 1: The overview of PNSGD unlearning. (Left) Proof sketch for PNSGD unlearning guarantees. (Right) PNSGD (un)learning processes on adjacent datasets. Given a mini-batch sequence $\mathcal{B}$, the learning process $\mathcal{M}$ induces a regular polyhedron where each vertex corresponds to a stationary distribution $\nu_{\mathcal{D}|\mathcal{B}}$ for each dataset $\mathcal{D}$. $\nu_{\mathcal{D}|\mathcal{B}}$ and $\nu_{\mathcal{D|\mathcal{B}}^{\prime}}$ are adjacent if $\mathcal{D},\mathcal{D}^{\prime}$ differ in one data point. We provide an upper bound $Z_{\mathcal{B}}$ for the infinite Wasserstein distance $W_{\infty}(\nu_{\mathcal{D}|\mathcal{B}},\nu_{\mathcal{D|\mathcal{B}}^{\prime}})$, which is crucial for non-vacuous unlearning guarantees. Results of [16] allow us to convert the initial $W_{\infty}$ bound to Rényi difference bound $d_{\alpha}^{\mathcal{B}}$, and apply joint convexity of KL divergence to obtain the final privacy loss $\varepsilon$, which also take the randomness of $\mathcal{B}$ into account.
@@ 1.1 Related Works
Machine unlearning with privacy guarantees. The concept of approximate unlearning uses a probabilistic definition of $(\epsilon,\delta)$-unlearning motivated by differential privacy [12], which is studied by [7, 8, 10]. Notably, the unlearning approach of these works involved Hessian inverse computation which can be computationally prohibitive in practice for high dimensional problems. Ullah et al. [5] focus on exact unlearning via a sophisticated version of noisy SGD. Their analysis is based on total variation stability which is not directly applicable to approximate unlearning settings and different from our analysis focusing on Rényi divergence. Neel et al. [9] leverage full-batch PGD for (un)learning and achieve approximate unlearning by publishing the final parameters with additive noise. Chien et al. [11] utilize full-batch PNGD for approximate unlearning with the analysis of Langevin dynamics. The adaptive unlearning requests setting is studied in [17, 6, 18], where the unlearning request may depend on the previous (un)learning results. It is possible to show that our framework is also capable of this adaptive setting since we do not keep any non-private internal states, though we only focus on non-adaptive settings in this work. We left a rigorous discussion for this as future work.
@@ 2 Preliminaries
Definition 2.2 (Rényi Differential Privacy (RDP) [19]).
@@
It is known to the literature that an $(\alpha,\varepsilon)$-RDP guarantee can be converted to the popular $(\epsilon,\delta)$-DP guarantee [12] relatively tight [19]. As a result, we will focus on establishing results with respect to $\alpha$ Rényi difference (and equivalently $\alpha$ Rényi difference). Next, we introduce our formal definition of unlearning based on $\alpha$ Rényi difference as well.
@@
Consider a randomized learning algorithm $\mathcal{M}:\mathcal{X}^{n}\mapsto\mathbb{R}^{d}$ and a randomized unlearning algorithm $\mathcal{U}:\mathcal{R}^{d}\times\mathcal{X}^{n}\times\mathcal{X}^{n}\mapsto\mathcal{R}^{d}$. We say $(\mathcal{M},\mathcal{U})$ achieves $(\alpha,\varepsilon)$-RU if for any $\alpha>1$ and any adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$, the $\alpha$ Rényi difference $d_{\alpha}(\rho,\nu^{\prime})\leq\varepsilon,$ where $\mathcal{U}(\mathcal{M}(\mathcal{D}),\mathcal{D}^{\prime})\sim\rho$ and $\mathcal{M}(\mathcal{D}^{\prime})\sim\nu^{\prime}$.
Our Definition 2.3 can be converted to the standard $(\epsilon,\delta)$-unlearning definition defined in [7, 8, 9], similar to RDP to DP conversion (see Appendix N). Since we work with the replacement definition of dataset adjacency, to unlearn a data point $\mathbf{d}_{i}$ we can simply replace it with any data point $\mathbf{d}_{i}^{\prime}\in\mathcal{X}$ for the updated dataset $\mathcal{D}^{\prime}$ in practice. One may also repeat the entire analysis with the objective function being the summation of individual loss for the standard add/remove notion of dataset adjacency. Finally, we will also leverage the infinite Wasserstein distance in our analysis.
@@
Algorithm 1 (Un)learning with PNSGD
1:  Parameters: stepsize $\eta$, noise standard deviation $\sigma$, dataset size $n$, mini-batch size $b$.
@@
2:  Minibatch Generation: randomly partition indices $[n]$ into $n/b$ mini-batches of size $b$: $\mathcal{B}^{0},\ldots,\mathcal{B}^{n/b-1}$.
@@
3:  Learning process $\mathcal{M}(\mathcal{D})$: given a dataset $\mathcal{D}=\{\mathbf{d}_{i}\}_{i=1}^{n}\in\mathcal{X}^{n}$, sample initial parameter $x_{0}^{0}$ from a given initialization distribution $\nu_{0}$ supported on $\mathcal{C}_{R}$. The output is the last iterate $x_{T}^{0}$.
4:  for  epoch $t=0,\ldots,T-1$ do
5:     for  iteration $j=0,\ldots,n/b-1$ do
@@
6:        $x^{j+1}_{t}=\Pi_{\mathcal{C}_{R}}\left(x^{j}_{t}-\eta g(x^{j}_{t},\mathcal{B}^{j})+\sqrt{2\eta\sigma^{2}}W^{j}_{t}\right),$ where $g(x^{j}_{t},\mathcal{B}^{j})=\frac{1}{b}\sum_{i\in\mathcal{B}^{j}}\nabla f(x^{j}_{t};\mathbf{d}_{i})$ and $W^{j}_{t}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(0,I_{d})$.
7:     end for
8:     $x^{0}_{t+1}=x^{n/b}_{t}$
9:  end for
@@
10:  Unlearning process $\mathcal{U}(\mathcal{M}(\mathcal{D}),\mathcal{D}^{\prime})$: given an updated dataset $\mathcal{D}^{\prime}=\{\mathbf{d}_{i}^{\prime}\}_{i=1}^{n}\in\mathcal{X}^{n}$ and current parameter $y_{0}^{0}$, the output is the last iterate $y_{K}^{0}$.
11:  for  epoch $k=0,\ldots,K-1$ do
12:     for  iteration $j=0,\ldots,n/b-1$ do
@@
13:        $y^{j+1}_{k}=\Pi_{\mathcal{C}_{R}}\left(y^{j}_{k}-\eta g(y^{j}_{k},\mathcal{B}^{j})+\sqrt{2\eta\sigma^{2}}W^{j}_{k}\right),$ where $g(y^{j}_{k},\mathcal{B}^{j})=\frac{1}{b}\sum_{i\in\mathcal{B}^{j}}\nabla f(y^{j}_{k};\mathbf{d}_{i}^{\prime})$ and $W^{j}_{k}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(0,I_{d})$.
14:     end for
15:     $y^{0}_{k+1}=y^{n/b}_{k}$
16:  end for
2.1 Converting Initial $W_{\infty}$ Distance to Final Rényi Divergence Bound
An important component of our analysis is to leverage the result of [16], which is based on the celebrated privacy amplification by iteration analysis originally proposed in  [20] and also utilized for DP guarantees of PNSGD in [15]. The goal of [16] is to analyze the mixing time of the PNSGD process, which can be viewed as the contractive noisy iterations due to the contractiveness of the gradient update under strong convexity assumption.